OpenStudy (anonymous):

A direct variation function contains the points (2, 14) and (4, 28). Which equation represents the function?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@Bossvideogamer21

OpenStudy (bossvideogamer21):

yeah i be here

OpenStudy (bossvideogamer21):

give me a minute to solve this

OpenStudy (anonymous):

okay @Bossvideogamer21

OpenStudy (bossvideogamer21):

The general way the books tell you to solve these, is to use the direct variation thing y = k*x and use your points (x,y) in there to figure out what k is.

OpenStudy (bossvideogamer21):

(x,y) = (2,14), x=2 and y=14 Put that into y = k*x

OpenStudy (bossvideogamer21):

That is the basic way to solve the direct variation problems, use y = k*x put a given point in, and find out what 'k' is

OpenStudy (bossvideogamer21):

holler if u need me